determine the equation of the parabola with x-intercepts -4 and 3, and that passes through (2,7)
Answer by stanbon(75887)

determine the equation of the parabola with x-intercepts -4 and 3, and that passes through (2,7)
-----------------
If -4 and 3 are x-intercepts, f(x) has factors (x+4) and (x-3)
-------------------
So, y = a(x+4)(x-3)
If the parabola passes thru (2,7)
7 = a(2+4)(2-3)
7 = a*6*-1
a = -7/6
------------------
EQUATION:
y = (-7/6)(x+4)(x-3)
=====================
Cheers,
Stan H.